Let P(T) be the probability of visiting a physical therapist and P(C) be the probability of visiting a chiropractor.
We know:
Using the formula for the union of two events:
P(T or C) = P(T) + P(C) - P(T and C)
Since P(neither T nor C) = 0.08, we have:
P(T or C) = 1 - 0.08 = 0.92
Substitute the known values:
0.92 = P(T) + P(C) - 0.28
0.92 = (P(C) + 0.16) + P(C) - 0.28
0.92 = 2P(C) + 0.16 - 0.28
0.92 = 2P(C) - 0.12
1.04 = 2P(C)
P(C) = 0.52
Therefore, P(T) = P(C) + 0.16 = 0.52 + 0.16 = 0.68
The probability of a randomly selected person visiting a physical therapist is 0.68 or 68%.
Snap a photo of your homework and get a step-by-step answer in seconds. You can ask follow-up questions until it makes sense.
Solve my question free Math help Find a math tutor© 2026 SnapAndSolve